Find the number of find the number of moles of oxygen required to convert 6.60 DM cube at SO2 measured at STP to SO3 molar volume of any gas at STP equal to 22.4 DM cube Avogadro's number 6.02 * 10 raise to power 23
1
Expert's answer
2020-05-21T14:54:22-0400

1mole= 6.02*10^23 avagadros number =22.4Dm

"2SO2 + O2=2SO3"

2SO2 =6.60DM

O2= 3.30DM

If one mole is 22.4DM , then 3.30DM is

(3.30/22.4)*1mole =0.147moles